Rarig Companies has decided to ask suppliers to bid on the 150,000 cartons of machine screws

that Rarig needs to purchase per year to support its manufacturing needs over the next five years.

You have decided to submit a bid to supply the machine screws. It will cost you $780,000 to

install the necessary equipment which will be fully depreciated over a straight line basis over five

years assuming zero salvage value. You estimate that you can sell the equipment for $50,000 at

the end of the project. Fixed costs will be $240,000 a year and variable costs will be $8.50 per

carton. You will also need to $75,000 in raw material inventory to begin the project and NWC

will stay constant over the five years (no dollars will need to be spent towards NWC in the other

years to support the project). At the end of year 5, you will sell the $75,000 worth of raw material

inventory you will have. You have a 16% cost of capital and a 35% tax rate. What is the

minimum bid price per carton that you should bid?
Answer #1

CF0=-Cost of equipment-working capital
CF1, CF2, CF3, CF4, CF5=((number of cartons*(price per carton-variable cost per carton)-fixed costs per year-depreciation)*(1-tax rate)+depreciation)
Additional cash flow in year 5=Salvage value*(1-tax rate)+working capital
Depreciation=Initial cost of equipment/5
NPV=CF0+CF1/(1+r)+CF2/(1+r)^2+CF3/(1+r)^3+CF4/(1+r)^4+CF5/(1+r)^5+Additional cash flow in year 5/(1+r)^5

At breakeven, NPV=0
-780000+50000*(1-35%)/1.16^5-75000+75000/1.16^5+((150000*(P-8.50)-240000-780000/5)*(1-35%)+780000/5)/16%*(1-1/1.16^5)=0
=>P=12.057882